OpenStudy (anonymous):

Given the equation y − 3 = (x + 6) in point-slope form, identify the equation of the same line in slope-intercept form. x − 2y = −12 y = x y = x + 6 y = x + 9

OpenStudy (anonymous):

slope intercept form is y = mx + b

OpenStudy (anonymous):

by solving for Y you get y = x + 6 + 3 y = x + 9

OpenStudy (anonymous):

Those it work that way for any equation similar to this? oh and Thank You.

OpenStudy (anonymous):

Yes that is the formula so it will work for all problems alike. Basically if you know the formula then you are just rearranging to get there.

OpenStudy (anonymous):

and you are welcome.

OpenStudy (anonymous):

Great !